OpenStudy (anonymous):

(x + 4 / 2x + 6)+(3 / x^2 - 9) Rational expressions, adding and subtracting help por favor.(x + 4 / 2x + 6)+(3 / x^2 - 9) Rational expressions, adding and subtracting help por favor.@Mathematics

OpenStudy (anonymous):

x+4 + 3 ---- ---------- 2(x+3) (x+3)(x-3) (x+4)(x-3)+3(2) x^2+x-6 (x+3)(x-2) x-2 ---------------- = --------- = ---------- = -------- 2(x+3)(x-3) 2(x+3)(x-3) 2(x+3)(x-3) 2(x-3)

OpenStudy (hoblos):

x+4 3 = -------- + --------- 2(x+3) (x+3)(x-3) (x+4)(x-3) +6 = ---------------- 2(x+3)(x-3) x^2 +x -12 +6 = -------------- 2(x+3)(x-3) (x+3)(x-2) = ---------- 2(x+3)(x-3) x-2 = ------- 2x-6